Electron-Muon Data for Run 1A and 1B

18

Jet Multiplicity

Following plot illustrates the serious S/N problem

• Event distribution versus minimum jet multiplicity • Require jet ET > 15 GeV (corrected)

b Quark Tagging in W+Jet Events

Applied b tagging techniques to W+jet events

• The W + 1 and W + 2 jet samples are ultimate control – Don’t expect significant top contribution

– Theory predicts low b hadron content

Sample  SVX bkg SVX tags  SLT bkg  SLT tags
W+1 jet  50 ± 12  40  159 ± 25  163
W+2 jet  21 ± 7  34  46 ± 7  55
W+≥ 3 jet  6.7 ± 2.1  27  15.4 ± 2.3  23